### v1

@@ Abstract
Context: Disasters are a common global occurrence with climate change leading to increase both their frequency and intensity. To reduce the impact of these disasters on lives and livelihoods it is important to provide accurate warnings and information about recovery and mitigation. Today most emergency management agencies deliver this information via mobile apps.
Objective: There are a large collection of disaster mobile apps available across the globe. But a detailed study is not yet conducted on these apps and their reviews to understand their key features and user feedback. In this paper we present a comprehensive analysis to address this research gap.
Method: We conducted a detailed analysis of 45 disaster apps and 28,161 reviews on these apps. We manually analysed the features of these 45 apps and for review analysis employed topic modelling and sentiment analysis techniques.
Results: We identified 13 key features in these apps and categorised them in to the 4 stages of disaster life cycle. Our analysis revealed 22 topics with highest discussions being on apps alert functionality, app satisfaction and use of maps. Sentiment analysis of reviews showed that while 22% of users provided positive feedback, 9.5% were negative and 6.8% were neutral. It also showed that signup/signin issues, network issues and app configuration issues were the most frustrating to users. These impacted user safety as these prevented them from accessing the app when it mattered most.
Conclusions: We provide a set of practical recommendations for future disaster app developers. Our findings will help emergency agencies develop better disaster apps by ensuring key features are supported in their apps, by understanding commonly discussed user issues. This will help to improve the disaster app eco-system and lead to more user friendly and supportive disaster support apps in the future.
@@
mobile apps, disaster apps, crisis management, user reviews, analysis, recommendations
@@ 1 Introduction
Natural disasters – such as floods, heatwaves, earthquakes, and bushfires – have become increasingly common. Driven by climate change, these disasters are affecting regions that previously experienced minimal impact, such as the floods in Oman in April 2024 [1]. These impacts are anticipated to only escalate in the future. Such disasters lead to loss of lives and livelihoods, disrupt education, harm health and hinder economic growth.
@@
However, many communities can reduce the impact of these disasters by improving their resilience with better preparation via information on better disaster management strategies before, during and after disasters. Traditionally, such information was communicated via broadcast media such as television, radios or via words of mouth. However, with advances in technology, software has started taking a prominent place in providing such information. With 85% of the today’s global population being mobile phone users, increasingly this information is provided via mobile apps [2, 3].
@@
Disaster apps are more and more considered to be critical support applications due to their ability to impact life and death situations [4, 5]. Currently most disaster apps are built and maintained by emergency agencies. These apps tend to have a localized focus on a specific country or a state. This is due to the inherently localised nature of most disasters and the need to incorporate localised tacit knowledge such as evacuation routes, into these apps. Therefore most local emergency agencies seek to develop their own version of a disaster app version. This has led to large number of disaster apps available in the Google Play Store and Apple Store. However, several issues have been reported in these apps, such as complicated interfaces, missed alerts and overwhelming information load [6, 4, 3]. In acute situations such as disaster led emergencies, even seemingly minor issues can become critical concerns that can lead to a decision between life and death.
Consider an example of a poorly designed disaster app. Firstly, due to bugs in the app, it may issue delayed bush fire alerts or miss alerts completely, which can lead to loss of lives and livelihoods. Secondly, during cascading disasters such as floods followed by landslides, due to poor color choices for each disaster, it may be difficult to locate impacted areas due to burst of colors in the map interface. Thirdly, due to issues in privacy and security the app may get hacked which can lead to issuing incorrect alerts that leads to massive community panics. Finally, it may lack critical advice on escape from and key support after the disaster, leading to lack of key information for users. Therefore, it is important to develop these apps methodically and by following good software engineering principles, ensuring key disaster support features are built and to high quality. When developing these apps, rather than re-inventing the wheel and each local emergency agency developing a disaster app from scratch, they can benefit by referring to existing apps. However, to the best of authors knowledge, a study analysing the wide range of disaster app features and their reviews on both Google Play Store and the App Store is currently lacking. In this study we aim to address this research gap.
@@
In this paper, we present a detailed analysis of all disaster apps that are available on Google Play Store and App Store and provide a list of commonly used features across all apps. We hope this will provide emergency agencies a pool of features to choose for their own app. We also analyse user reviews for these apps via topic modelling techniques to identify topics that are commonly discussed by users. We then conduct a sentiment analysis on these reviews to understand common preferences and pain points of the users. This analysis will help agencies to understand what aspects of the app should receive more attention and what is important for users.
@@ 2 Related Work
2.1 Mobile Applications for Disaster Management
Disaster apps have a significant diversity in the type of information they provide and their intended target groups. The type of information vary based on type of disaster and the stage of disaster life cycle they focus on [5]. The type of disaster can be a single disaster such as floods or may present information about multiple disasters in which case they are known as multi-hazard apps. The stages of a disaster life cycle are four fold according to disaster management literature. These are ’Preparation’, ’Response’, ’Recovery’ and ’Mitigation’ [7]. The disaster apps can focus on one or many of these stages. The preparation stage involves information such as capacity building, monitoring and providing early warnings; the response stage involves information for evacuation, search and rescue; the recovery stage involves damage assessment and reconstruction; and lastly the mitigation stage involves information such as risk analysis and risk appraisals [8]. The target groups can be the global community for international apps and local communities such as state, regional communities for local apps.
Disaster apps are broadly classified into two categories: general purpose apps and built for purpose apps. General purpose apps include three sub-categories One-to-one, One-to-many and Many-to-many [9]. One-to-one refers to messaging apps such as Whatsapp and Wechat which are used by a person to send personal message to another. One-to-many refers to news apps are maintained by news agencies such as CNN and while publishing other news, they also publish disaster news to the public. Many-to-many include social media apps such as Twitter and Facebook. Out of these, people tend to favour familiar platforms such as social media which they have frequently used before the disaster occurrence [10].
However, disaster management authorities have concerns in promoting the use of general-purpose platforms for emergency situations, as many issues arise such as privacy, information quantity, and content quality [9]. Built for purpose apps are usually maintained by emergency agencies and will focus on functions such as Alert and information dissemination: disseminating authorised information before and during disasters; and Information collation: gathering, filtering and analysing data to build situation awareness. These apps are usually one-way; commonly originating from the authorities to the public. Examples of such apps are the American Red Cross Apps and the Federal Emergency Management Agency (FEMA) App, both of which were developed by their respective agencies.
@@ 2.2 App Review Analysis
Given that app reviews include rich information, a growing number of researchers and practitioners have investigated app reviews to extract actionable insights for different software engineering tasks [11, 12]. App reviews have been mainly analyzed to support requirements engineering tasks. For example, the studies [13, 14] developed automated approaches to identify new features requested by users in mobile app reviews. Some have extracted non-functional requirements from app reviews (e.g., [15, 16]. Several studies have focused on identifying issues (e.g., UI issues, crashes, bugs) that users face when working with apps by mining user reviews (e.g., [17, 18, 19, 20]. Another line of research attempted to (automatically) detect fake reviews and non-informative reviews (e.g., [21, 22]). Another important research theme is the sentiment analysis of app reviews to detect users’ emotions about the app (e.g., [23, 24]). Finally, some research has recently focused on manually or automatically identifying and classifying human, social, and ethical factors (e.g., human and ethical values [25, 26, 27, 28] and privacy [29]) from app reviews.
2.3 Disaster Mobile App Analysis
@@
Researchers have sought to explore disaster mobile apps in a limited number of studies. Tan et al. [9] systematically reviewed crisis informatics literature, including disaster apps. But these focused on apps discussed in academic literature, as we found in our analysis; not all of them were available to the public as some were still in the proof of concept level [9]. Corcuera et al. [30] also looked into disaster apps. Still, their literature review was limited to apps in the Google Play store, and they sought only to explore the functionalities of these apps and evaluate their effectiveness for disaster risk reduction [30]. They did not explore the user reviews that are available in these apps for insights into user perception of features and app performance.
@@
Studies analyzing app reviews, described in Section 2.2, typically do not cover disaster apps. This is mainly because these studies choose the reviews from the most popular apps (e.g., Instagram and TikTok apps) and app categories (e.g., health, social, communication) or randomly select reviews in Google Play Store and App Store. The popularity of an app is usually defined by the number of downloads, which excludes disaster apps from consideration. This exclusion also applies to other less frequently downloaded but important app categories, such as depression apps [28] and sports apps [31]. However, as explained earlier, disaster apps are used in critical situations where even a minor issue can make a difference between life and death. Therefore, a detailed study of disaster apps available to the public today is long overdue.
Figure 1: Overview of our study approach
@@ 3.1 Research Questions
Rationale. We want to explore the perceptions and opinions of disaster app users that are expressed through the reviews of these mobile apps. By employing topic modeling techniques, this RQ extracts and categorizes the primary topics discussed in user reviews. This analysis, covering 28,161 reviews across 45 mobile apps, identified 22 distinct topics.
@@
Rationale. We want to develop deeper understanding of key grievances and dissatisfaction among users of disaster mobile apps. This was analysed with the ’user ratings’ provided for each app, and via a sentiment analysis performed on the reviews.
@@ 3.2 Data Collection
Figure 1 summarizes our research methodology. This section details the steps involved in identifying, gathering, and preparing the data to answer our RQs. The specific methodologies employed to address each research question are outlined in Section 4.
@@ 3.2.1 A Dataset of Disaster Mobile Apps
Step 1: App Collection This step involved locating apps for disasters. We used three approaches to achieve this.
Approach 1: Literature-based app collection: We first extracted a list of mobile apps used for disaster management via the systematic review of Tan et al., which identified 29 mobile apps [9]. We extracted 5 apps from the Apple Store and Google Play Store. The rest of the apps discussed in this paper were still in the research stage and, hence, were not published at the Apple Store or Google Play Store. They thus lack user reviews, finalised features and real end users.
Approach 2: Keywords-based app collection. Due to the limited apps that were available via the literature-based app collection approach, we searched for relevant apps in both Google and Apple app stores in different parts of the world. We identified disaster apps using disaster-specific keywords [32], such as ‘flood’, ‘earthquake’, and ‘tornado’. These keywords were used to search in the titles and descriptions of apps in the Play Store and Apple Store. This process led to an identification of 38 apps tailored to addressing natural disaster situations.
@@
To address this challenge we adopted search approach 3 that involved conducting a google search for potential disaster support apps in each country. We used the query “Emergency apps in [Country]” as emergency apps often included functionalities related to natural disasters. We did not use “Disaster apps in [Country]” as a query because this would be overly restrictive and may not yield a complete app list. Some of the disaster apps obtained through this method include ‘Yurekuru Call’ and ‘Safety Tips’. These apps were found in a blog written by Robin Lewis [33], which was discovered using the query “Emergency apps in Japan.” This approach was repeated for all 196 countries worldwide [34].
@@
With approach 3, we found 73 disaster-related apps developed by government and non-government entities, available on the Google Play Store and Apple App Store. However, this was much less than what we expected for an overall of 200 countries in the world. While it was understandable that some developing countries may not have an app due to resource limitations, we did not feel that this justified all the missing countries. Therefore, to understand why many countries haven’t developed specific disaster apps, we visited official state government websites to find how they provided disaster information to their citizens. Our research showed that each country had various methods of disseminating disaster information, many often without relying on mobile apps. For instance, Austria utilizes sirens  [^1] https://www.bmi.gv.at/204_english/skkm/warning.aspx, Canada informs their citizens via television and radio  [^2] https://www.alertready.ca, Saudi Arabia sends notifications directly to personal devices  [^3] https://998.gov.sa/English/Pages/Automatic-Early-Warning.aspx, India employs SMS  [^4] https://sachet.ndma.gov.in, and Belgium opts for email  [^5] https://www.be-alert.be/en/how-does-be-alert-work. These are only one of the methods that are used by these countries as most countries prefer to employ multiple methods for disseminating disaster information.
@@
Step 2: Removing Duplicates. Figure 1 shows that 116 apps were obtained from our three app collection methods. This included 5 apps from the literature review, 38 apps from keyword-based searches, and 73 apps from country-based searches (See Figure 1). However, there were overlaps in apps that were collected from some methods. For instance, the ‘FEMA’ app, designed for disaster preparedness in the USA, was identified across all three methods. Similarly, the literature-based review and the keyword search found the ‘Disaster Alert’ app, which provides notifications for disasters in the Pacific region. Once these duplicates were removed we had a collection of 99 apps.
Step 3: Removing False Positives. From an initial collection of 99 apps, a filtering process was undertaken to remove false positive apps [31]. False positives refer to apps that do not have features pertinent to natural disasters. Such general-purpose emergency apps tended to focus on providing services to connect with police or fire, provide assistance during medical emergencies, traffic incidents, and other non-disaster-related emergencies. Most of these apps were added to our data set as a result of our Approach 3 in data collection. To assess the 99 apps accurately, we examined their descriptions and downloaded them to directly inspect the available features. This evaluation identified 45 apps that have natural disaster-related functionalities as shown in Figure 1. The list of apps can be found in Appendix Tables 10, 11and 12.

### v2

@@ Abstract
Disaster support mobile apps play a central role during emergencies in conveying timely alerts to communities and providing recovery guidance. Their importance continues to grow as climate change increases the frequency and intensity of disasters, and as mobile devices become a primary information source for most people. However, despite their critical role, there has been limited systematic investigation into how these apps are designed and how users experience them in real-world contexts. This paper presents a large-scale empirical study of disaster support mobile apps aimed at addressing this gap. We investigated 44 apps available on the Google Play Store and the App Store and analyzed 28,161 user reviews to understand common functionalities, user expectations, and challenges encountered during use. Our feature analysis identifies 13 core capabilities spanning the stages of preparation, response, and recovery of the disaster lifecycle. Topic Modelling revealed 18 topics, with the highest discussions focusing on app alert functionality, the use of maps, and monetary transactions. Sentiment analysis revealed that while 53.09% of users provided positive feedback, 24.30% expressed negative opinions, and 11.25% remained neutral. It also showed that signup/signin issues, network problems, and app crashes were the most frustrating to users. These impacted user safety by preventing access when it mattered most. Our findings emphasize the need to approach disaster apps as safety-critical and human-centred systems, where reliability and ease of use are essential. We provide practical recommendations to support developers and emergency agencies in improving the dependability and accessibility of future disaster support apps.
@@
disaster support mobile apps, safety-critical software, feature analysis, human-centered software
@@ 1 Introduction
Natural disasters, including floods, heat waves, earthquakes, and wildfires, have become increasingly common. According to the Atlas of Mortality and Economic Losses from Weather [UNFCCC2021], natural disasters have increased fivefold over that 50-year period. Such disasters lead to loss of life and livelihood, disrupt education, harm health and hinder economic growth.
@@
However, many communities can reduce the impact of these disasters by improving their resilience through better preparation and access to information on disaster management strategies before, during, and after disasters. Traditionally, such information was shared through broadcast media such as television, radio, or word of mouth, and later through short message service (SMS) alerts. While SMS can deliver brief warnings, mobile apps now provide more advanced, interactive, and adaptable platforms. With 85% of today’s global population using mobile phones, disaster-related information is increasingly delivered through mobile apps [HowManyP98_online, tan2020usability].
@@
Disaster applications are increasingly regarded as critical support tools due to their potential to influence life or death situations [tan2020understanding, dallo2021should]. Consequently, many individual developers, emergency agencies, and government organizations have begun developing and publishing their own versions of disaster apps on the Google Play Store and the App Store. However, several issues have been reported in these apps, such as complicated interfaces, missed alerts and overwhelming information load [CFAappgl73_online, tan2020understanding, tan2020usability]. In acute situations such as disaster-related emergencies, even seemingly minor issues can become critical concerns that can lead to a decision between life and death.
Consider an example of a poorly designed disaster app. Firstly, due to bugs in the app, it may issue delayed bush fire alerts or completely miss alerts, which can lead to loss of lives and livelihoods [TheGuardian2025]. Secondly, poorly designed map visualizations, such as rainbow color palettes, can mislead users. Studies show that non-uniform color scales in flood hazard maps distort how people judge severity and extent of risks [Jenny2024]. Thirdly, due to privacy and security issues, the app may get hacked which can lead to issuing incorrect alerts that leads to massive community panics [Heckel2020]. Therefore, it is important to develop these apps methodically and by following good software engineering principles, and to ensure that key disaster support features are built to a high standard.
When developing these apps, rather than re-inventing the wheel and each local emergency agency developing a disaster app from scratch, they can benefit from referring to existing apps. Moreover, since disasters often share similar patterns across regions, the findings from one country can inform improvements in others. However, to the best of authors’ knowledge, a study analyzing the wide range of disaster app features and their reviews on both the Google Play Store and the App Store is currently lacking. In this study we aim to address this research gap.
@@
We present a detailed analysis of disaster-related mobile apps available on the Google Play Store and the App Store, and provide a comprehensive list of commonly implemented features across these apps. This work aims to offer emergency agencies a feature pool that can guide the development of their own applications. In addition, we analyze user reviews using topic modeling techniques to identify the key themes discussed by users. A sentiment analysis is then performed to uncover common preferences and pain points, providing insights into the aspects of these apps that require greater attention.
@@ 2 Related Work
2.1 Mobile Apps for Disaster Management
Disaster apps have a significant diversity in the type of information they provide and their intended target groups. The type of information vary based on type of disaster and the stage of disaster lifecycle they focus on [dallo2021should]. The type of disaster can be a single disaster such as floods or may present information about multiple disasters in which case they are known as multi-hazard apps. The stages of a disaster lifecycle are fourfold according to disaster management literature. These are ’Preparation’, ’Response’, ’Recovery’ and ’Mitigation’ [04Unit4d19_online]. The disaster apps can focus on one or many of these stages. The preparation stage involves information such as capacity building, monitoring and providing early warnings; the response stage involves information for evacuation, search and rescue; the recovery stage involves damage assessment and reconstruction; and lastly the mitigation stage involves information such as risk analysis and risk appraisals [poser2010volunteered]. The target groups can be the global community for international apps and local communities such as state, regional communities for local apps.
Disaster apps are broadly classified into two categories: general purpose apps and purpose-built apps. General purpose apps include three sub-categories One-to-one, One-to-many and Many-to-many [Tan2017]. One-to-one refers to messaging apps such as Whatsapp and Wechat which are used by a person to send personal message to another. One-to-many refers to news apps are maintained by news agencies such as CNN and while publishing other news, they also publish disaster news to the public. Many-to-many include social media apps such as Twitter and Facebook. Out of these, people tend to favor familiar platforms such as social media which they have frequently used before the disaster occurrence [HADDOW201453].
However, disaster management authorities have concerns in promoting the use of general-purpose platforms for emergency situations, as many issues arise such as privacy, information quantity, and content quality [Tan2017]. purpose-built apps are usually maintained by emergency agencies and will focus on functions such as Alert and information dissemination: disseminating authorized information before and during disasters; and Information collation: gathering, filtering and analyzing data to build situation awareness. These apps are usually one-way; commonly originating from the authorities to the public. Examples of such apps are the NINA-Die Warn-App des BBK and the Federal Emergency Management Agency (FEMA) App, both of which were developed by their respective agencies.
@@ 2.2 App Review Analysis
Given that app reviews include rich information, a growing number of researchers and practitioners have investigated app reviews to extract actionable insights for different software engineering tasks [dkabrowski2022analysing, genc2017systematic]. App reviews have been mainly analyzed to support requirements engineering tasks. For example, one study [Tan2020Modified] collected and thematically analyzed online reviews of disaster apps. From these insights, the authors built a conceptual usability framework specific to disaster apps. Researchers [johann2017safe, de2021re, Motger2024] developed approaches to identify new features requested by users in mobile app reviews. Some have extracted non-functional requirements from app reviews (e.g., [jha2019mining, lu2017automatic]). Other work has focused on detecting key features to determine which aspects users value most or complain about the most [Wu2021]. Additional efforts have aimed at identifying issues (e.g., UI issues, crashes, bugs) that users face when working with apps by mining user reviews (e.g., [man2016experience, chen2021should, maalej2015bug, mcilroy2016analyzing, Motger2024]). Another line of research attempted to (automatically) detect fake reviews and non-informative reviews (e.g., [martens2019towards, noei2019too]). Another important research theme is the sentiment analysis of app reviews to detect users’ emotions about the app (e.g., [martens2017emotion, luiz2018feature]). Finally, some research has recently focused on identifying and classifying human, social, and ethical factors (e.g., human and ethical values [shahin2023study, obie2023automated, shams2020society, bowie2022user] and privacy [nema2022analyzing]) from app reviews.
2.3 Review-based Disaster Mobile App Analysis
@@
Researchers have explored disaster mobile apps in a limited number of studies. Tan et al. [Tan2017] systematically reviewed crisis informatics literature, including disaster apps. However, these studies mainly focused on apps discussed in academic literature, and as we found in our analysis, not all of them were available to the public. Some were still at the proof-of-concept stage and some can’t be found in any website and application store. Corcuera et al. [corcuera_2022] also looked into disaster apps. Still, their literature review was limited to apps in the Google Play Store, and they sought only to explore the functionalities of these apps and evaluate their effectiveness for disaster risk reduction [corcuera_2022]. They did not explore the user reviews that are available in these apps for insights into user perception of features and app performance.
@@
Studies analyzing app reviews, described in Section 2.2, typically do not cover disaster apps. This is mainly because these studies choose the reviews from the most popular apps (e.g., Instagram and TikTok apps) and app categories (e.g., health, social, communication) or randomly select reviews in the Google Play Store and the App Store. The popularity of an app is usually defined by the number of downloads, which excludes disaster apps from consideration. This exclusion also applies to other less frequently downloaded but important app categories, such as depression apps [bowie2022user] and sports apps [Chembakottu2023]. However, as explained earlier, disaster apps are used in critical situations where even a minor issue can make a difference between life and death. Therefore, a detailed study of disaster apps available to the public today is long overdue.
@@ 3.1 Research Questions
Rationale. We want to explore the perceptions and opinions of disaster app users that are expressed through the reviews of these mobile apps. By employing topic modeling techniques, this RQ extracts and categorizes the primary topics discussed in user reviews. This analysis, covering 28,161 reviews across 44 mobile apps, identified 18 distinct topics.
@@
Rationale. We want to develop deeper understanding of key grievances and dissatisfaction among users of disaster mobile apps. This was analyzed with the ’user ratings’ provided for each app, and via a sentiment analysis performed on the reviews.
Figure 1: Overview of our study approach
@@ 3.2 Data Collection
Figure 1 summarizes our research methodology. This section details the steps involved in identifying, gathering, and preparing the data to answer our RQs.
@@ 3.2.1 A Dataset of Disaster Mobile Apps
Step 1: App Collection. This step involved locating apps for disasters. We used three approaches to achieve this. The collection process was conducted between 19 October and 25 October 2023. Applications published outside this timeframe were not included in our study.
Approach 1: Literature-based app collection. We first reviewed existing literature to identify studies that compiled lists of disaster management applications. From the works in [Tan2017, Tan2020Modified], we obtained a combined list of 115 mobile apps related to disaster management. At this stage, duplicate entries were not removed, as the deduplication process was performed later after integrating this list with the apps collected through the other two approaches.
Approach 2: Keywords-based app collection. Due to the limited apps that were available via the literature-based app collection approach, we searched for relevant apps in both the Google Play Store and the App Store in different parts of the world. We identified disaster apps using disaster-specific keywords [Navarro2021], such as ‘flood’, ‘earthquake’, and ‘tornado’. These keywords were used to search in the titles and descriptions of apps in the Google Play Store and Apple Store. This process led to an identification of 54 apps tailored to addressing natural disaster situations.
@@
To address this challenge we adopted search approach 3 that involved conducting a google search for potential disaster support apps in each country. We used the query “Emergency apps in [Country]” as emergency apps often included functionalities related to natural disasters. We did not use “Disaster apps in [Country]” as a query because this would be overly restrictive and may not yield a complete app list. Some of the disaster apps obtained through this method include ‘Yurekuru Call’ and ‘Safety Tips’. These apps were found in a blog written by Robin Lewis [lewis:2018], which was discovered using the query “Emergency apps in Japan.” This approach was repeated for all 196 countries worldwide [britannica].
@@
With approach 3, we found 73 disaster-related apps developed by government and non-government entities, available on the Google Play Store and the App Store. However, this was much less than what we expected for an overall of 196 countries in the world. While it was understandable that some developing countries may not have an app due to resource limitations, we did not feel that this justified all the missing countries. Therefore, to understand why many countries have not developed specific disaster apps, we exhaustively reviewed every official state government website to identify how they provided disaster information to their citizens. Our research showed that each country had various methods of disseminating disaster information, many often without relying on mobile apps. For instance, Austria utilizes sirens  [^1] https://www.bmi.gv.at/204_english/skkm/warning.aspx, Canada informs their citizens via television and radio  [^2] https://www.alertready.ca, Saudi Arabia sends notifications directly to personal devices  [^3] https://998.gov.sa/English/Pages/Automatic-Early-Warning.aspx, India employs SMS  [^4] https://sachet.ndma.gov.in, and Belgium opts for email  [^5] https://www.be-alert.be/en/how-does-be-alert-work. These are only one of the methods that are used by these countries as most countries prefer to employ multiple methods for disseminating disaster information.
@@
Step 2: Combining List of Apps and Removing the Duplicates. Figure 1 shows that 242 apps were obtained from our three app collection methods. This included 115 apps from the literature review, 54 apps from keyword-based searches, and 73 apps from country-based searches. However, there were overlaps in apps that were collected from some methods, thus we remove the duplicates.
Step 3: Filtering Apps. After removing duplicates we get 190 apps. To refine this collection, we implemented a two-stage filtering process. First, we excluded applications that were not available in either the Google Play Store or the App Store, since the absence of availability prevented us from retrieving user reviews. Second, we eliminated false positives [Chembakottu2023]. False positives were defined as applications that did not include functionalities relevant to natural disasters. Specifically, we excluded applications that did not provide at least one feature listed in Table 1. For instance, some applications were designed to facilitate communication with law enforcement authorities. While these address emergency situations, they fall outside the scope of our study, which is limited to natural disaster contexts. Similarly, applications providing assistance in medical emergencies (e.g., cardiac arrest) were excluded because their focus is on general health crises rather than disaster-specific scenarios. This rationale also applied to generic social media and crowdfunding platforms, which, although they may contain incidental information about disasters, are not explicitly developed for disaster management. To ensure the accuracy of this second filtering stage, we downloaded selected applications to verify their functionalities directly. Following this, we identified 44 applications that incorporated features directly related to natural disaster management (See Table 1). The complete list of applications detailed in Appendix Table A2 and Table LABEL:tbl:list_of_play_store_app.